### v1

# UDEEP: Edge-based Computer Vision for In-Situ Underwater Crayfish and Plastic Detection

Abstract
Invasive signal crayfish have a detrimental impact on ecosystems. They spread the fungal-type crayfish plague disease (Aphanomyces astaci) that is lethal to the native white clawed crayfish, the only native crayfish species in Britain. Invasive signal crayfish extensively burrow, causing habitat destruction, erosion of river banks and adverse changes in water quality, while also competing with native species for resources and leading to declines in native populations. Moreover, pollution exacerbates the vulnerability of White-clawed crayfish, with their populations declining by over 90% in certain English counties, making them highly susceptible to extinction. To safeguard aquatic ecosystems, it is imperative to address the challenges posed by invasive species and discarded plastics in the United Kingdom’s river ecosystem’s. This article introduces the UDEEP, Underwater Deep learning EdgE computing Platform, for the detection of crayfish and plastic. The paper also presents two publicly available underwater datasets, annotated with sequences of crayfish and underwater plastic debris, captured in freshwater rivers under varying visibility conditions. The crayfish and plastic object detection training achieved the best accuracy of 0.9 mean Average Precision (mAP) @0.5 Intersection over Union (IoU) on You Only Look Once (YOLO) v5s. YOLO v8n achieved the highest precision at 92%. Subsequently, YOLO v5n achieved the best recall score overall at 93%. As for the amount current required for inference, YOLO v8s required the least amount of current for inference, 2548mA and YOLO v5n consumed the highest current at 3277mA.
The UDEEP platform can play a crucial role in environmental monitoring by performing on-the-fly classification of Signal crayfish and plastic debris while leveraging the efficacy of Artificial Intelligence (AI), Internet of Things (IOT) devices and the power of edge computing (i.e., NVIDIA Jetson Nano (NJN)). By providing accurate data on the presence, spread and abundance of these species, the UDEEP platform can contribute to monitoring efforts and aid in mitigating the spread of invasive species.
keywords:
Deep Learning (DL), Signal crayfish, Invasive Species, Plastics, image segmentation, YOLO v5, YOLO v8
[…]
[NTU]organization=Nottingham Trent University,Department of Computer Science addressline=Clifton Campus, city=Nottingham, postcode=NG11 8NS, country=UK
[…]
[MAI]organization=Mediprospects AI, addressline=5-7 High Street, city=London, postcode=E13 0AD, country=UK
{graphicalabstract}
Research highlights {highlights}
UDEEP was developed for automated crayfish and plastic monitoring.
YOLO v5s,5n,8s and 8n models were trained and evaluated on an NJN board.
The models accurately detected crayfish and plastics in images under varying underwater conditions.
Two publicly available underwater datasets are introduced, annotated with sequences of Signal crayfish and plastic debris.
1 Introduction
Invasive species exert substantial impacts on ecosystems globally, comparable to the effects of pollutants such as discarded plastic. Extensive information exists about these species, including their characteristics, invaded regions, invasion pathways, and impacts Tekman et al. (2022); Hobbs (2000). However, there is still a need for comprehensive assessments that consider multiple aspects of the invasion process simultaneously Bernery et al. (2022); Woods et al. (2016).
Invasive signal crayfish (Pacifastacus Leniusculus) pose a threat to native species and habitats, causing substantial economic consequences. These invasive crayfish out-compete and prey on native white-clawed crayfish (Austropotamobius Pallipes), causing population declines and even extinctions Mathers et al. (2016). The white-clawed crayfish is a keystone species found in freshwater bodies in Western Europe Mirimin et al. (2022). However, its populations have experienced significant declines due to the presence of a contagious fungal–type disease (Aphanomyces Astaci) originating from North America, competition with non-indigenous crayfish species, and habitat deterioration Mirimin et al. (2022). Additionally, their burrowing activities result in habitat degradation, destabilising riverbanks and disrupting aquatic vegetation. The economic impact of these damages includes the loss of biodiversity, reduced productivity in fisheries, and the need for habitat restoration, which can require substantial financial investments Faller et al. (2016). Access to accurate and comprehensive data on these invasions is crucial Bernery et al. (2022).
In the process of detecting and recognising aquatic animals, the initial step involves acquiring data and gathering information. However, due to the complexity and dynamicist of the aquatic ecosystem, collecting underwater images can be relatively difficult and even dangerous Li et al. (2023). Nevertheless, advancements in technology have made it possible to collect fine-scale data using unmanned systems and a variety of sensors, such as physical and environmental data, imagery, and bathymetry data. Bathymetry is the measurement of depth of water in rivers, lakes or oceans to produce depth maps. These can then be used to characterise details of animal behaviour and their interactions within the ecosystem Ling et al. (2016). DL is also being used to identify and locate aquatic species in raw images, videos, and audio recordings for bio-monitoring purposes Goodwin et al. (2022).
Use of cameras for biodiversity monitoring has less impact on biodiversity and can cover larger areas and for longer periods compared to traditional observation methods, that involve observers walking in transects. Data that can be collected includes abundance, density, species richness and proportion of area occupied Stephenson (2020). Additionally, Light Detection and Ranging (LIDAR) detection can be added to measure ranges, temperature, light intensity sensors can be added to provide a more holistic dataset.
This paper introduces the Underwater UDEEP which combines various techniques, including background subtraction for selective cropping of regions of interest in high-resolution images for the detection of crayfish and underwater plastic debris. The platform plays a pivotal role in detecting and estimating the density of Signal crayfish and plastic debris by leveraging the efficacy of AI, IOT devices and the power of edge computing (i.e., NJN). By delivering precise data regarding the presence, distribution, and population density of these species and discarded plastics, the UDEEP platform significantly contributes to monitoring endeavours and facilitates efforts to mitigate the proliferation of invasive species and environmental pollution by plastic debris.
The remainder of this article includes the literature review in Section 2, the description of the computational experiments for training, testing and validation with YOLO variants 5s and 5n, 8s and 8n in Section 3, the results of the computational experiments Section 4, and finally the conclusions and future work in Section 5
2 Related Work
Recent years have seen a significant increase in the interest in processing underwater images Mittal et al. (2022). The interest is driven by the need to study and monitor aquatic plants and animals, which has applications in marine biology, economy, and biodiversity management. Understanding the behaviour and numbers of aquatic species is crucial for various purposes, including protecting endangered species, analysing differences in species, and early detection of climatic events like pollution and global warming. For instance, Plankton, which produces over 80% of the world’s oxygen Witman and Witman (2021), plays a critical role in the ocean’s food chain and atmosphere-ocean connection. Monitoring plankton levels is essential, as both low and excessive levels can have harmful consequences. Furthermore, organisms like Posidonia Oceanic, which thrive in clean water, are vital for biodiversity, beach erosion reduction, and water quality enhancement Mittal et al. (2022). Image processing complements other techniques like physiochemical water analysis and sonar-based detection in understanding the impact of global warming and human activity on marine life.
2.1 The challenges of underwater monitoring
Underwater imaging contends with substantial energy loss, resulting in diminished and inconsistent illumination and visibility, particularly in the deeper expanses Mittal et al. (2022). The situation is further exacerbated by the presence of freshwater and ocean currents, intensifying the degradation of visibility. Impurities, suspended solids, low contrast, and compromised edges and details further impact underwater images. The distortion of colour based on distance, caused by non-uniform spectral propagation, adds another layer of complexity. Additionally, underwater conditions present intricate and uncontrolled backgrounds, posing challenges in segregating subjects such as sea cucumbers and shellfish from their surroundings. The unbridled movement of fishes in three-dimensional space, often concealed behind objects, compounds the difficulty in accurately determining their pose, orientation, and size. The scarcity of large-scale, publicly available datasets for underwater images, coupled with imbalances, necessitates the incorporation of data augmentation techniques.
The limited availability of training data complicates the training of deep Convolutional Neural Network (CNN). Moreover, the severe constraints on wired/wireless communication underwater make streaming of underwater imaging exceptionally challenging, resulting in lower-resolution images. This limitation hampers the distinction of features, diminishing the effectiveness of conventional techniques. Conventional Computer Vision (CV) methods lean on hand-crafted features like Scale-Invariant Feature Transform (SIFT), Histogram of Oriented Gradients (HOG), and local binary patterns. However, these features lack generalisation capabilities across diverse classes, scenarios, and datasets. The accuracy of traditional techniques plateaus with increasing training data, constraining their overall performance. Extracting hand-crafted features demands domain expertise and time, rendering them less practical for underwater image processing. Previous work on fish recognition often involves dead fish or fish in unnatural conditions, such as swimming pools, limiting its relevance to live marine environments.
2.2 Early advances in underwater bio-diversity monitoring
In the early 90’s, Strachan et al. (1990) laid the groundwork for automated fish grading based on size and colour, utilising image analysis to estimate fish shape and colour. This innovative approach allowed for the differentiation of haddock fish stocks from the Rockall Plateau, located midway between Greenland and Iceland and the North Sea. The method involved digitised Red, Green, and Blue (RGB) images captured from beneath a conveyor belt, processed using bespoke software written in the C programming language.
The study marked a significant advancement in distinguishing between fish species. Building on the pioneering work, Storbeck and Daan (2001) conducted further research in 2001 by integrating computer vision with a neural network, expanding on the foundations set by Strachan et al. (1990).
Both studies required the fish to be laid flat on a conveyor belt moving at a specific speed, with a vertical arrangement of a laser source and camera above the belt. However, the classification method, while groundbreaking, proves impractical for underwater environments.
2.3 Modern application of computer vision for underwater bio-diversity monitoring
A more contemporary perspective is presented in the work of Zhang et al. Zhang et al. (2021), showcasing a modern application of computer vision for underwater object detection. Employing the YOLO v4 model, the authors achieved an mAP of 92.65% at a processing speed of approximately 45 Frames per Second (FPS). Despite the success, it’s crucial to note that the achievement relied on a high-end NVIDIA graphics card for accelerating the YOLO v4. However, deploying such hardware alongside sensors in physical underwater settings becomes unfeasible due to size and power constraints.
In response to the constraint, it is advisable to explore embedded systems as viable alternatives. Such platforms present a more pragmatic deployment option for sensors, ensuring the practical implementation of advanced computer vision techniques in underwater environments. As underwater computer vision emerges as a burgeoning field, studying analogous works becomes imperative to identify potential gaps and opportunities that can be leveraged to advance knowledge in this domain.
For instance, Hegde et al. (2021) successfully implemented a trained MobileNetV2 model capable of running on a Raspberry Pi, achieving an impressive precision of 98% for mackerel. While acknowledging that a YOLO model could offer faster results, they deemed it unfeasible due to hardware limitations. Consequently, their image capture and processing occurred at a rate of once every 2 seconds, equivalent to about 0.5 FPS. This limitation could prove critical in scenarios involving Underwater Autonomous Vehicles, where objects of interest might traverse the sensors before processing completion.
The incorporation of a YOLO v5 model on the NJN platform is set to enhance performance, delivering a finely tuned solution tailored for the target applications. Noteworthy is the study by Hegde et al. (2021), which had a limited implementation of comprehensive image pre-processing methods, mainly focusing on cropping. In contrast, our UDEEP algorithm surpasses these limitations by integrating background subtraction which is a technique enables to extract moving objects (foreground) from the background, as discussed by Huang et al. (2017) and Kalsotra and Arora (2022).
The sophisticated enhancement is anticipated to streamline object identification, potentially fortifying our algorithm’s efficacy in detecting small objects Hegde et al. (2021). Importantly, our dataset boasts an extensive collection of [450 500] photos per each category, significantly contributing to the robustness of our approach. In a comparable study, Wang et al. (2020) employed YOLO nano on an NJN to accelerate the algorithm at the edge. The model used by Wang et al. Wang et al. (2020) demonstrated the capability to detect underwater objects at an impressive frame rate of 8 FPS, accompanied by a notable mAP of approximately 74.8%. Notably, a discernible trade-off emerges between runtime and accuracy when juxtaposed with the work of Hegde et al. (2021). The strategic use of image pre-processing might provide an avenue for simpler, faster models to attain higher accuracy in underwater object detection scenarios.
Chen et al. (2021) presented a methodology similar to the one employed in this work, utilising data augmentation to enhance the capabilities of the object detection model. Wei et al. (2020) proposed mosaic augmentation as use to augment the dataset to enhance the detection of smaller objects in cluttered environments. Moreover, Wei et al. Wei et al. (2020) introduced an innovative adaptation, involving the calculation of gains on the RGB channels to modify and adjust images, accommodating variations in underwater light levels. The UDEEP distinguishes itself through the deliberate exclusion of the background. This intentional omission is crafted to emphasise the salient features within the image, establishing our methodology as distinct from approaches that encompass the entire scene.
Considerable efforts have been directed towards advancing efficient autonomous underwater Underwater Autonomous Vehicles (UAVs) employing machine learning algorithms for debris detection. Leveraging the computational power of an NVIDIA Jetson TX2 Graphics Processing Unit (GPU), Fulton et al. (2019) conducted comprehensive experiments involving transfer learning on various models, including YOLO v2, Tiny-YOLO, Faster R-CNN, and Single Shot Detector (SSD). In their approach, the author fine-tuned each network individually for R-CNN, SSD, and Tiny-YOLO. However, for YOLO v2, an additional step involved training using transfer learning. The models were specifically tailored for underwater plastic debris detection, and the results indicated that Faster R-CNN achieved the highest accuracy, albeit with longer inference times. YOLO v2 struck a balance between speed and accuracy, SSD exhibited the best inference time, and Tiny-YOLO emerged as the top performer in terms of overall performance.
Saleh et al. (2024) assert that automated fish phenotyping, encompassing parameters such as fish length, weight, size, and mortality, provides invaluable insights into understanding the comprehensive profile of classified fish. The task of individually counting fish is notably resource-intensive, demanding skilled labour for hours, thereby contributing to elevated project costs and extended delivery timelines. In a related domain, Ranjan et al. (2023) introduced a real-time mortality monitoring system based on YOLO v7, achieving a commendable mAP of 93.4% and an F1 score of 0.89. The UDEEP builds upon this foundation, not only incorporating these crucial findings, but also extending the phenotyping metrics. Specifically, we introduce age estimation, distinguishing between adult and juvenile crayfish.
3 Methodology
The Methodology section is divided into four parts. The first part introduces the UDEEP platform workflow and followed by a review of the hardware setup. The third section explains the YOLO v5n, YOLO v5s, YOLO v8n and YOLO v8s model architectures and the final section reviews the dataset and delves into the data augmentation techniques applied to the crayfish and underwater plastic datasets.
3.1 UDEEP Platform Workflow
The UDEEP platform workflow involves two distinct steps. Initially, we categorise the dataset into different sets based on the objects of interest, crayfish or plastic. Subsequently, we train the system using CNN with these grouped images, encompassing various objects of interest. In the second step, during testing, we employ the Gaussian Mixture Model (GMM) to segment the testing images. Each segment was tested to determine if it corresponds to any of the specified objects of interest. Figure 1 below illustrates the proposed UDEEP workflow.
Figure 1: The proposed UDEEP platform workflow.
3.1.1 Training
The training phase is the initial step of any classification algorithm. This is undertaken on a given set of images that have been selected due to objects of interest being present in them. For the UDEEP platform, the images contained crayfish and plastic debris in the underwater environment. This phase is further divided into two sub-phases namely Segregation and feature generation.
3.1.2 Segregation
The images used in this Research project were sourced from rivers within the United Kingdom and contain crayfish and contain crayfish and plastics. CNNs don’t perform well in extracting features from complex scenes, they however excel in learning features when training images showcase single objects with minimal noise. However, in multi-object images, CNNs attempt to learn features collectively, making it challenging to precisely describe individual objects. The bounding box technique was employed to delineate and select the object(s) of interest in each image via the Roboflow platform Roboflow (2023). A bounding box serves as the minimum enclosing box around a specific object, capturing the entire object of interest while minimising noise. All training images are resized to a $416\times 416$ pixels.
3.1.3 Testing
To assess an underwater image, it’s essential to identify the various objects within it. The authors employed the GMM segmentation to determine the number of objects in an image, the classifications, and their outline. Subsequently, these segmented objects are fed into the trained CNN model to validate whether a given object is of interest or not.
3.1.4 segmentation
The task of image segmentation can be expressed as categorising pixels into semantic labels (semantic segmentation), delineating individual objects (instance segmentation), or a combination of both (panoptic segmentation). Semantic segmentation involves labelling pixels at the pixel level with specific object categories (such as crayfish or plastic) for all pixels in the image. Minaee et al. (2021)
GMM-based models, which consider pixel intensities, are employed for improving the robustness to Gaussian noise. The GMM method categorises an input image matrix into segments using the complete data log-likelihood concept. Each pixel’s Gaussian density function, multiplied by the probability of belonging to a segment, yields the log-likelihood based on GMM-generated segments. Similar to other segmentation methods, GMM faces the challenge of selecting the number of segments, commonly addressed by analysing the image histogram.
3.2 UDEEP underwater monitoring prototype: AI inference at the edge
A meticulously designed and implemented prototype for underwater monitoring is illustrated in Figure 2. The system features an NJN equipped with an Intel RealSense D435i depth sensor. The NJN is used for accelerating the YOLO v5 and YOLO v8 inference at the edge. The NJN selected is tailored for diverse tasks, including image identification, object detection, data segmentation, and speech processing. It encompasses a robust configuration, boasting a 128-core Maxwell GPU, a quad-core ARM A57 Application Processor Unit (APU) running at 1.43 GHz, and 4 GB of 64-bit LPDDR4 memory operating at a speed of 25.6 GB/s. This system is further enhanced by the inclusion of an NJN, preloaded with a pre-trained DL model.
Figure 2: UDEEP monitoring system.
3.3 Network Architecture
Four object classification network architectures were used from the YOLO family used in this research project, driven by the necessity to perform object detection at the edge where processing and power resources are constrained. The choice leaned towards the compact and efficient small and nano versions of YOLO.
YOLO v5n is a lightweight version of the YOLO v5 object classification algorithm, Ultralytics (2021). It was the first classification DL classification architecture and the smallest and fastest model in the YOLO v5 family due to striking a balance between model weight and detection accuracy. It has improved detection speed while maintaining a high degree of accuracy compared to previous YOLO versions. Liu et al. (2023).
YOLO v5s achieves a balance between detection accuracy and speed, stood out as another significant choice. The object classification algorithm contains fewer layers and parameters compared to other YOLO v5 models, YOLO v5s prioritised efficiency and ease of training, making it suitable for real-time applications. Despite potential sacrifices in detection accuracy compared to larger models, YOLO v5s is used in various domains, including liquid biopsy of lung cancer and pouring robot object detection Zhao et al. (2023).
Moving to the YOLO v8n, the smallest and fastest model in the YOLO v8 family with 3.2 million parameters, it introduced several architectural improvements over YOLO v5 Selcuk and Serif (2023). YOLO v8n has improved accuracy, faster object detection speed, higher mAP, and the exclusion of anchor boxes during object detection. The enhanced speed and accuracy made YOLO v8n particularly well-suited for real-time detection tasks, aligning seamlessly with the project’s objectives Selcuk and Serif (2023); Wang et al. (2023).
The fourth model, YOLO v8s has more parameters than the other three YOLO models used in this research, 11.2 million parameters—8 million more than YOLO v8n. This increased size should translate to improved performance. The model has more Floating Point Operations Per Second (FLOPS) than its smaller YOLO v8n counterpart Selcuk and Şerif (2023); Selcuk and Serif (2023); Wang et al. (2023).
3.4 Datasets Pre-processing
The authors have previously generated a public custom underwater crayfish dataset, consisting of 2486 images Machado et al. (2023). This dataset was partitioned into three subsets: 1740 images for the training set, 249 for testing, and 497 images for validation. Additionally, another public custom dataset, also created by the authors, focuses on plastic and comprises 1220 images Machado (2022). Similarly, the plastic dataset was divided into three segments: 854 images for the training set, 122 for testing, and 244 images for validation. A sample is shown in Figure 3.
Figure 3: Sample images from datasets used in training DL models for detecting crayfish and plastic in UK rivers.
The dataset contains images captured from real underwater environments that contain crayfish and a wide variety of plastic debris in different stages of decay and are either clearly visible or covered with varying amounts of overgrowth or other underwater debris. Furthermore, lighting quality and water clarity vary greatly across the dataset which makes the images resemble real-world conditions that the UDEEP prototype is intended to work in.
Image labelling is an important step whereby labels or tags are assigned to images based on their content in order to categorise and organise them in a machine-readable format. The Roboflow computer vision platform, Roboflow (2023) was used to accomplish the labelling task. Using Roboflow, the dataset was labelled with two classes, defined as Crayfish: all signal crayfish or Plastic: all plastic debris/material.
3.5 Dataset Augmentation
Fine-tuning YOLO models for optimal performance goes beyond merely implementing the algorithm itself. An essential aspect of the labelling process is the utilisation of image augmentation techniques, such as cropping, flipping, rotating, and colour shifting, which can enhance the model’s generalisation and robustness Xu and Matzner (2018). Data augmentation, which is not properly customised, may not yield desired model performance. This section delves into various data augmentation techniques performed on the crayfish and underwater plastic datasets introduced in this article.
  • 1.
    Image Scale augmentation:
    Image Scale augmentation is the action of resizing the input images into various dimensions or scales in order to train the YOLO model on a dataset containing a variety of scales. This adapts the model to objects of differing sizes, which is the case in real-world scenes. The images in the dataset were stretched to 416x416
  • 2.
    Image Flip: Horizontal, Vertical: Images in the dataset were vertically and horizontally flipped. Vertically flipping creates a mirror image where the top becomes the bottom and vice versa. This helps the YOLO model detect objects that may appear upside down in real-world scenarios. Similarly, flip left-right augmentation horizontally flips an image, allowing the model to learn and detect objects from different perspectives. Training on these flipped images enhances the model’s robustness, making it adaptable to accurately detect objects regardless of their orientation
  • 3.
    Cropping: In cases where the subject is obscured by other objects, only a portion of the object can be captured in the photograph. Therefore, we randomly cropped the images, with both the cropping width and height ranging from 0% Minimum Zoom, to 49% Maximum Zoom.
  • 4.
    Mosaic Augmentation: Mosaic augmentation was also applied to the dataset. This is a method that combines multiple images into a single mosaic-like training sample. This technique enhances the YOLO model’s ability to detect objects in complex scenes with overlapping or crowded environments. By training with mosaic-augmented images, the model becomes more adept at handling situations where objects are partially hidden or blend together, ultimately improving its accuracy in challenging scenarios.
  • 5.
    Hue, Saturation, Exposure and Gray-scale: The hue component of images in the dataset was altered in the range between -25° and +25°, saturation altered in the range between -42% and +42%, exposure altered between -22% and +22% and Gray-scale applied to 47% of images.
    Adjusting the hue component mimics diverse lighting conditions, including daylight or artificial lighting. This allows the model to train and recognise objects effectively under a range of illumination settings Applying the Hue, Saturation and exposure augmentation ensures the trained model is robust enough to handle real world lighting exposure, various colour schemes and contrasts
4 Result Analysis
This section presents the results for training, testing and validation. In the previous section, a classifier was trained with 2300 images in the training set and evaluated its performance on 72 images in the test set. The classifier’s accuracy was assessed based on three measures: training accuracy, test accuracy, and validation accuracy. Training accuracy refers to how accurately the classifier performed on the training set, while test accuracy measures its performance on the test set. Validation accuracy indicates the rate at which the classifier successfully classified newly encountered images in the validation set. Additionally, we considered the recorded losses to evaluate the effectiveness of our approach.
The models will be evaluated according to the highest mAP, F1 score and average precision. The mAP, weighted mean of precision’s at each threshold, (i.e a comparison score between the ground-truth bounding box to the actual detected box), is useful where many classes are present in object detection with each class having differing detection performance. A higher mAP score indicates a higher detection accuracy. The formulae for Average Precision (AP) is rulled by Equation 1 below. mAP is the mean of AP over all the queries.
[…]
where $R_{n}$ and $P_{n}$ are the precision and recall at the $n$th threshold.
Two functions that return an array of labels was created to produce 18 different confusion matrices. The first function grabs all the true values from the test set. The second function grabs all the predicted labels from each model with the highest mAP value per augmentation. The two arrays of values are then used to plot the confusion matrices. The confusion matrix uses recall equation, precision, and F1-score, equations 2, 3 and 4 respectively.
$Recall:re=\frac{tp}{tp+fn}$ (2)
$Precision:pr=\frac{tp}{tp+fp}$ (3)
$F1-score:f1=2\times\frac{pr\times re}{pr+re}$ (4)
where, $tp$ is the true positives, $tn$ is the true negatives, $fp$ is the false positives and $fn$ is the false negatives.
4.1 Experimental Results on the Crayfish and Plastic Dataset
In this article, the YOLO v5n, YOLO v5s, YOLO v8n and YOLO v8s model architectures are trained on the Underwater plastic and crayfish datasets. After training, the four models are applied to the test set of the Underwater plastic and crayfish datasets and measured results of crayfish and plastic detection are obtained.
The figure 4 below present the precision-recall curves scores obtained during the classification task. The graphs all show a high area under the curve, which signifies a combination of high recall and high precision. High precision indicates that the model achieved a low false positive rate, meaning that the classifier accurately identifies crayfish without many false alarms. On the other hand, high recall corresponds to a low false negative rate, implying that the classifier successfully captures the majority of positive results.
When both precision and recall scores are high, it indicates that the classifier is returning accurate and relevant results (high precision) while also capturing a significant portion of all positive results (high recall).
Consequently, an analysis of the results of our training indicate that YOLO v8n achieved the highest precision among the four YOLO architectures achieving a score of 0.92 followed by YOLO v8s that achieved a score of 0.893. YOLO v5s achieved a score of 0.885 and finally YOLO v5n achieving an F1 score of 0.85. These results are shown in the graphs below.
Figure 4: YOLO v5n,YOLO v5s,YOLO v8n and YOLO v8n PR curve.
Figures 5 below presents the F1 confidence score results for the crayfish and plastic classification task. YOLO v5s and YOLO v8s achieved the highest scores at F1 score of 87%. This was followed by YOLO v8n that had an F1 score of 85% and finally YOLO v5s achieving the lowest score of 84%. A summary of these results is also presented in table 1.
Figure 5: YOLO v5n,YOLO v5s,YOLO v8n and YOLO v8n F1-Score.
To measure current usage, Jetson-stats, Bonghi (2019) package was used to analyse the average current used by the board while running the classification models. A python script was adapted to utilise the Jetson-stats library to record the current measurements into a Comma Separated Values (CSV) file stored on the board. The mean inference current consumed during the classification task on the NVIDIA Jetson Nano was calculated based on data collected from classifying 69 images. Measuring power consumption is crucial to guarantee that the device’s power requirements are adequately addressed. Effective power source planning is essential, as it enables the implementation of backup measures for devices operating remotely at the edge. Such proactive measures aim to minimise the potential risk of downtime caused by power outages.
Table 1: Yolo training results.
Class Precision Recall F1 Score Mean Inference Current (mA)
YOLO v5s 0.885 0.92 0.87 3158
YOLO v5n 0.851 0.93 0.84 3277
YOLO v8s 0.917 0.92 0.85 2548
YOLO v8n 0.83 0.92 0.87 2710
Table 1 shows the combined validation results for training both the crayfish and plastics and the mean inference current in mA consumed by the NJN board during inference. The results show that YOLO v8s requires the least amount of current and YOLO v5n requires the most current for inference.
4.2 UDEEP In-Situ underwater Testing
To showcase the effectiveness and benefits of the UDEEP underwater object detection approach employing the four YOLO network models, experiments were conducted in real underwater environments in Rivers within the East of England. The experiment involved fully submerging the UDEEP prototype in the river where crayfish sightings had been reported.
Appropriate permission from the landowners, Health and safety training and licensing had to be obtained. The UDEEP prototype operated fully off-grid due to its electric power source being supplied by rechargeable batteries charged by solar power cells. Figure 6 displays visualisations of the detected objects as part of the experiment results.
Figure 6: UDEEP crayfish and plastic detection results.
5 Conclusions and Future Work
This article presents the Underwater DL edge computing platform (UDEEP), a novel monitoring platform for underwater crayfish and plastics identification that is world-leading in terms of originality, significance and rigour. The platform integrates diverse methods, including background subtraction for selectively cropping regions of interest in high-resolution images to detect crayfish and underwater plastic debris. By providing accurate data on the presence, distribution, and population density of these species, the UDEEP platform makes a substantial contribution to monitoring initiatives and aids in addressing the spread of invasive species A public custom underwater crayfish dataset, consisting of 2486 images Machado et al. (2023) was generated as part of this Research Project and made publicly available. Additionally, a second public custom dataset, also created by the authors, focuses on plastic and comprises 1220 images Machado (2022) was generated and released into the public domain.
This paper further presents experiments using YOLO v5 nano and small and YOLO v8 nano and small image classification algorithms that were trained to identify crayfish and plastic in underwater environments using a NJN board. From the results, it is evident that the two versions of YOLO were able to detect and count the number of crayfish and plastic with a high degree of accuracy. The results of measuring the inference current showed that YOLO v8s required the least amount of current and YOLO v5n required the most amount of current among the four models.
Future work involves integrating the models into other versions of edge computing devices for example the NVIDIA TX2 which will significantly enhance the generation of statistics related to the detection, quantity, and location of invasive and endangered species. The UDEEP platform will empower policymakers, conservation authorities, and researchers with real-time data on the spread of invasive non-native species and discarded plastic debris floating along water bodies. The availability of such data will facilitate the enactment of well-informed conservation laws and the implementation of effective mitigation measures to safeguard native endangered species and reduce plastic pollution.
While the current study concentrated on identifying discarded plastic and signal crayfish as the non-native species responsible for the decimation of native white-clawed crayfish, the UDEEP platform has the potential to be further trained to recognise other non-native species. For instance, it could be extended to identify invasive non-native species such as the red-eared terrapin and Chinese mitten crab, broadening the scope of its applicability and environmental impact assessment.